OpenStudy (onthecly2):

Need urgent help with some trig! Solve the given equation over the interval [0, 2π): 2 cos x − 1 = 0.

OpenStudy (akshay18):

(pi/3,5pi/3)

OpenStudy (akshay18):

just draw cosx graph and see the intersection with y=1/2

OpenStudy (akshay18):

how to post image here

OpenStudy (irishboy123):

\(\implies \cos x = \dfrac{1}{2}\) no nice answers, so drawing is good idea The computer spits this out :( \(x = \pi \dfrac{1}{3} (6 n - 1)\) Draw is good

OpenStudy (akshay18):

to solve any trignimentry que you need to understand geaphs properly

OpenStudy (akshay18):

ask more que please

OpenStudy (irishboy123):

oh Scheisse wasn't paying attention :( missed fundie relationship.